# Doppler Correspondence: Non-Iterative Scan Matching With Doppler Velocity-Based Correspondence

### v1

@@ Abstract
Achieving successful scan matching is essential for LiDAR odometry. However, in challenging environments with adverse weather conditions or repetitive geometric patterns, LiDAR odometry performance is degraded due to incorrect scan matching. Recently, the emergence of frequency-modulated continuous wave 4D LiDAR and 4D radar technologies has provided the potential to address these unfavorable conditions. The term 4D refers to point cloud data characterized by range, azimuth, and elevation along with Doppler velocity. Although 4D data is available, most scan matching methods for 4D LiDAR and 4D radar still establish correspondence by repeatedly identifying the closest points between consecutive scans, overlooking the Doppler information. This paper introduces, for the first time, a simple Doppler velocity-based correspondence—Doppler Correspondence—that is invariant to translation and small rotation of the sensor, with its geometric and kinematic foundations. Extensive experiments demonstrate that the proposed method enables the direct matching of consecutive point clouds without an iterative process, making it computationally efficient. Additionally, it provides a more robust correspondence estimation in environments with repetitive geometric patterns.
@@ I Introduction
Recently, the frequency-modulated continuous wave (FMCW) 4D LiDAR and 4D radar have shown great potential in alleviating these limitations. Unlike the conventional LiDAR, FMCW-based 4D ranging sensors are more resilient to weather conditions [8, 13]. Moreover, they provide 4D point cloud information (range, azimuth, elevation, and Doppler velocity), which has potential [6, 13] for robust scan matching in scenarios with repetitive geometric structures.
@@ II Reltaed Work
4D LiDAR is an emerging technology that enables Doppler velocity measurements, offering advantages in environments with geometric degeneracy or dynamic objects. Several recent studies have explored the potential of this novel sensing modality for odometry and SLAM tasks. Wu et al. [18] introduced the first continuous-time 4D LiDAR-only odometry method leveraging Doppler velocity measurements. Their method employs Gaussian process regression to estimate vehicle trajectories and correct motion distortion caused by scanning, significantly outperforming existing methods. Similarly, Hexsel et al. [6] proposed DICP, which integrates Doppler velocity measurements into the ICP framework for robust point cloud registration. By jointly optimizing Doppler and geometric objective functions, their method demonstrated improved registration accuracy and convergence speed, particularly in featureless environments such as tunnels and hallways. To further enhance computational efficiency, Yoon et al. [20] developed a lightweight odometry method that directly estimates 6-DOF vehicle velocity from Doppler velocity without data association. By incorporating gyroscopic angular velocity measurements, their method achieves real-time performance with high computation efficiency, presenting a trade-off between accuracy and efficiency. Zhao et al. [24] proposed FMCW-LIO, a LiDAR-inertial odometry, that uses Doppler velocities for motion compensation and dynamic point removal, achieving robust state estimation in repetitive geometric environments. Pang et al. [14] proposed an efficient 4D LiDAR odometry framework. Using scan slicing and vehicle kinematics within an iterative error state kalman filter (IESKF), their method reduces computational costs while maintaining high accuracy.
@@ II-C Summary and Limitations of 4D Radar / LiDAR Odometry
The 4D information of FMCW 4D LiDAR and 4D radar has leveraged its advantages in three main directions. First, Doppler velocity has been effectively incorporated into optimization frameworks to improve accuracy and robustness in challenging environments, such as repetitive geometric settings, foggy conditions, or a significant number of dynamic objects [3, 4, 5, 6, 10, 11, 12, 13, 18, 22, 24]. Second, scan matching methods have been developed by refining point correspondences, using RCS information [7] or uncertainty of point clouds [19]. Third, lightweight and efficient odometry frameworks have focused on reducing computational costs while maintaining accuracy [14, 20].
While these advances have proven to be effective, most of the current correspondence between scans solely focuses on 3D geometric information, failing to fully leverage the Doppler velocity available in 4D data. Our work addresses this limitation by introducing a novel correspondence that explicitly incorporates Doppler information.
The remaining sections are organized as follows. In Section III, Doppler Correspondence is introduced in detail, including its geometric derivation and the rationale behind its robustness in repetitive geometric environments. Section IV presents the experimental results, demonstrating the effectiveness of the proposed method across various datasets and comparing its performance against point-to-point ICP and DICP. In Section V, the limitations of the approach are introduced, addressing scenarios where the method may face challenges. In Section VI, Doppler Correspondence is integrated with ICP to compare its performance with ICP and DICP, demonstrating its potential applicability. Finally, Section VII provides the conclusion, summarizing the key contributions of the work and its potential for the field of 4D ranging sensor-based odometry.
@@ III-A Problem Definition
    $(\delta R,\delta t)=\underset{\delta R,\delta t}{\arg\min}\sum_{(p,q)\in\mathcal{C}}\|q-(\delta Rp\;+\;\delta t)\|^{2}.$
@@ III-D Outlier Rejection
Any pair of points that significantly deviates from these constraints are considered outliers and are removed from the set of correspondences. By enforcing these thresholds, the impact of false correspondence estimation is mitigated.
@@
Figure 4: (a) illustrates multiple candidates due to symmetry in range and Doppler velocity and (b) shows that distance-based outlier rejection method filters out mismatched pairs.
@@ III-E Non-iterative Scan Matching Algorithm
The process begins by receiving two consecutive scans, $\mathcal{P}$ and $\mathcal{Q}$, and calculating $f(p_{i},v_{p,i})\text{ and }g(q_{j},v_{q,j})$. FindClosest is then performed to estimate correspondences. Unlike the conventional ICP explained in Algorithm 2, which relies on 3D-based correspondence, our method operates on 1D-based correspondence. This could reduce computational complexity, even though $f(p_{i},v_{p,i})$ and $g(q_{j},v_{q,j})$ need to be calculated. While this approach may be slightly less accurate than 3D-based methods due to less information, it offers faster computations. Furthermore, unlike the closest point correspondence methods [1, 6, 15, 16], which iteratively refines correspondences through multiple iterations, Algorithm 1 performs a direct correspondence estimation in a single step. This eliminates the iterative search process, significantly reducing the computational load, as illustrated in Fig. 1. Next section, the feasibility of Doppler Correspondence is validated by comparing the proposed method against ICP-based approaches.
@@
$\hat{R},\hat{t}\leftarrow\underset{R,t}{\arg\min}\sum_{(p,q)\in\mathcal{C}}\|q-(Rp+t)\|^{2}$ return $\hat{R},\hat{t}$
@@
      $\delta{R},\delta{t}\leftarrow\underset{R,t}{\arg\min}\sum_{(p,q)\in\mathcal{C}}\|q-(Rp+t)\|^{2}$
@@ IV-B Experiment Setup
The proposed method is evaluated by comparing its performance with established methods: point-to-point ICP and DICP [6]. DICP is a novel algorithm that integrates Doppler velocity into the Optimization in the ICP framework. It enhances scan matching by jointly optimizing both closest point term and additional constraints derived from the Doppler velocity of each point. These methods are tested in two conditions: No Seed and With Seed. In the No Seed category, no initial pose estimate is provided, whereas the With Seed category uses the estimated pose from the previous scan pair under a constant-velocity motion model assumption [17].
@@ IV-C Evaluation Metrics
To address the limitations of single-metric evaluations like the absolute trajectory error (ATE) [9], the performance is evaluated using the relative error (RE) from [23]. This metric standardizes trajectory estimation accuracy, allowing fair comparisons across different methods. The evaluation measures odometry accuracy through segment-wise error analysis. This approach provides detailed insights into the local and global performance of the trajectory estimation. By using shorter sub-trajectories, RE captures local consistency and short-term accuracy, while longer sub-trajectories emphasize global consistency and long-term accuracy. This dual capability makes RE a versatile and informative metric for trajectory analysis.
@@ IV-D Experiment Results
Figure 8: Visualization of established correspondences after a single iteration in sparse, semi-sparse, and dense point clouds.
@@ IV-F Correspondence Visualization
The matching results of Doppler Correspondence and the closest point correspondence method are visualized on sparse, semi-dense, and dense point clouds. The correspondences established after the single iteration is performed, as shown in Fig. 8. With only a single iteration, Doppler Correspondence achieves more coherent and consistent matches than the closest point approach across all levels of point cloud density.
@@ V Limitations
While Doppler Correspondence introduced in Section III offers notable advantages in terms of computational efficiency and robustness in repetitive geometric environments, it also comes with certain limitations. Specifically, the key scenarios where the method may fail are highlighted below. We also clarify the main assumptions underlying Eqs. (4) and (5). Furthermore, potential scenarios where these assumptions may fail in practice are discussed.
@@ V-A Key Assumptions and Their Breakdown
For identification of the case where the approximation in Eqs. (4) and  (5) holds, the motion of a point $p_{i}$ can be considered as a combination of translation $\Delta t\in\mathbb{R}^{3\times 1}$ and rotation $\Delta R\in SO(3)$ from the sensor between a time interval $\Delta T$. Suppose $p_{i}(T)$ is the point $p_{i}$ at time $T$. After the time interval $\Delta T$, the point moves according to:
@@
$\Delta R\;\approx\;I+\Delta T\,\omega^{\times}.$
@@
Derivation of Eq. (5) starts from the relationship $q_{j}=\Delta Rp_{i}+\Delta t$, which can be expressed as:
@@
$=-v_{q_{j}}+(\Delta T\omega^{\times}\rho)\cdot n_{q_{j}}.$
Rearranging them:
$\frac{q_{j}-p_{i}}{\Delta T}=\underbrace{v_{q_{j}}}_{\text{Doppler velocity term}}-\underbrace{(\Delta T\omega^{\times}\rho)\cdot n_{q_{j}}}_{\text{rotation term}}$ (14)
The difference between Eqs. (14) and (5) is the existence of rotation term. As $v_{q_{j}}=n_{q_{j}}\cdot\rho$, Eq. (14) can be written as:
$\frac{q_{j}-p_{i}}{\Delta T}=((I-\Delta Tw^{\times})\rho)\cdot n_{q_{j}}.$
If $\Delta T||w||_{2}\approx 0$, Eq. (14) becomes same as Eq. (5). Therefore, the sufficient condition of Doppler Correspondence is $\boldsymbol{\Delta T||w||_{2}\approx 0}$. Violating this condition may lead to degradation or failure of the proposed method. The following examples illustrate the cases where the assumption holds and it does not:
@@ V-A2 U-Turn (Assumption Fails)
V-B Strategies to Minimize Rotation-Induced Errors
The rotation-dominant situation of ground vehicles can be mitigated to some extent. In ground vehicles, the $\omega$ direction is similar to $z$-axis, and the $\rho$ direction is near the $x$-axis. Assuming $\omega=[0,0,0.56]^{T}$ and $\rho=[3,0,0]^{T}$, similar to U turn, the values of the rotation term (i.e., $(\Delta T\omega^{\times}\rho)\cdot n_{q_{j}}$ in Eq. (14)) can be expressed as a function of the azimuth and elevation of the normal vector $n_{q_{j}}$ as shown in Fig. 10. For the U turn example, if points with absolute azimuth angle less than 50 degrees and absolute elevation angle greater than 7.5 degrees are selected, the rotation error can be significantly reduced, as shown in Fig. 11. Despite the promising reduction in rotation error, more comprehensive investigations remain as future work.
Figure 10: Relationship between azimuth and elevation (in degree) of points on rotation term ($\Delta T\omega^{\times}\rho\cdot n_{q_{j}}$). The function decreases when the absolute value of azimuth is small and the absolute value of elevation is large.
Figure 11: Example of U-turn from the Loop1_4. (a) shows the estimated trajectory of the vehicle using Doppler Correspondence without any filtering method, while (b) shows the estimated trajectory with the filtering method applied.
VI Practical Use of Doppler Correspondence
Doppler Correspondence may offer versatile applications due to its simplicity. One straightforward implementation is its integration with ICP in a weighted form, as explained in Algorithm 3. The process begins by computing $f(\mathcal{P})$ and $g(\mathcal{Q})$ from the point clouds and establishing Doppler Correspondence. In this setup, correspondences estimated from Doppler Correspondence remain fixed throughout the ICP iterations, while closest point correspondences are dynamically updated. This combination enables the optimization process to leverage two distinct correspondence characteristics: spatial information and Doppler information.
To evaluate the effectiveness of the proposed approach, experiments are conducted using a semi-dense 4D radar dataset, Loop 1, and dense 4D LiDAR datasets, Straight Wall, Curved Wall, and Baker-Barry Tunnel (Empty). The Doppler Correspondence method is integrated into the ICP framework and compared against standard point-to-point ICP and DICP. The parameters are configured as follows: for the semi-sparse 4D radar dataset, $\tau_{\text{spatial}}=3$, $\tau_{\text{doppler}}=5$, and $\alpha=0.6$. For the dense 4D LiDAR datasets, the parameters are set to $\tau_{\text{spatial}}=0.5$, $\tau_{\text{doppler}}=5$, and $\alpha=0.9$.
The experimental results on the semi-dense 4D radar dataset in Table II and the dense 4D LiDAR dataset in Table III demonstrate that the proposed method outperforms point-to-point ICP and DICP in odometry performance. Notably, point-to-point ICP ($\alpha=0$) tends to diverge in repetitive geometric environments, whereas incorporating Doppler Correspondence helps to address this limitation effectively. By leveraging Doppler Correspondence, the approach successfully combines spatial and Doppler information, resulting in lower rotational and translational errors across different segment lengths.
Algorithm 3 Weighted Iterative Closest Point Algorithm Based on Doppler Correspondence
@@ VI Practical Use of Doppler Correspondence
      $\delta{R},\delta{t}\leftarrow\underset{R,t}{\operatorname{arg\,min}}\;(1-\alpha)\sum_{(p,q)\in\mathcal{C}_{\text{spatial}}}\|q-(Rp+t)\|^{2}+\alpha\sum_{(p,q)\in\mathcal{C}_{\text{doppler}}}\|q-(Rp+t)\|^{2}$
@@
TABLE II: Odometry performance on the semi-dense 4D radar dataset. Each cell reports the averaged rotational and translational errors (in degrees and meters) across various segment lengths. Bold values highlight the best performance for each segment.
Evaluation Units (rotation [deg] / translation [m])
Distance point-to-point ICP DICP Weighted ICP
8m 0.749 / 0.539 0.770 / 0.547 0.740 / 0.530
16m 1.159 / 0.958 1.203 / 0.976 1.133 / 0.941
24m 1.434 / 1.361 1.494 / 1.398 1.397 / 1.338
32m 1.686 / 1.748 1.760 / 1.807 1.639 / 1.721
40m 1.944 / 2.219 2.018 / 2.213 1.888 / 2.097
48m 2.187 / 2.500 2.251 / 2.613 2.121 / 2.466
TABLE III: Odometry performance on the dense 4D LiDAR dataset. The averaged rotational and translational errors (in degrees and meters) are presented for various segment lengths, with bold values indicating the best results for each segment. Note that Div. indicates the trajectory of the algorithm diverged.
Evaluation Units (rotation [deg] / translation [m])
Distance point-to-point ICP DICP Weighted ICP
8m Div. 1.028 / 0.289 0.964 / 0.259
16m Div. 1.650 / 0.514 1.498 / 0.454
24m Div. 2.512 / 0.779 2.042 / 0.663
32m Div. 3.313 / 0.963 2.366 / 0.829
40m Div. 3.964 / 1.140 2.882 / 1.015
48m Div. 4.783 / 1.498 3.079 / 1.402

### v2

@@ Abstract
Achieving successful scan matching is essential for LiDAR odometry. However, in challenging environments with adverse weather conditions or repetitive geometric patterns, LiDAR odometry performance is degraded due to incorrect scan matching. Recently, the emergence of frequency-modulated continuous wave 4D LiDAR and 4D radar technologies have provided the potential to address these unfavorable conditions. The term 4D refers to point cloud data characterized by range, azimuth, and elevation along with Doppler velocity. Although 4D data is available, most scan matching methods for 4D LiDAR and 4D radar still establish correspondence by repeatedly identifying the closest points between consecutive scans, overlooking the Doppler information. This paper introduces, for the first time, a simple Doppler velocity-based correspondence—Doppler Correspondence—that is invariant to translation and small rotation of the sensor, with its geometric and kinematic foundations. Extensive experiments demonstrate that the proposed method enables the direct matching of consecutive point clouds without an iterative process, making it computationally efficient. Additionally, it provides a more robust correspondence estimation in environments with repetitive geometric patterns. The implementation of our proposed method is publicly available at https://github.com/Tars0523/Doppler_Correspondence.
@@ I Introduction
Recently, the frequency-modulated continuous wave (FMCW) 4D LiDAR and 4D radar have shown great potential in alleviating these limitations. Unlike conventional LiDAR, FMCW-based 4D ranging sensors are more resilient to weather conditions [10, 15]. Moreover, they provide 4D point cloud information (range, azimuth, elevation, and Doppler velocity), which has potential [7, 15] for robust scan matching in scenarios with repetitive geometric structures.
@@ II Related Work
4D LiDAR is an emerging technology that enables Doppler velocity measurements, offering advantages in environments with geometric degeneracy or dynamic objects. Several recent studies have explored the potential of this novel sensing modality for odometry and SLAM tasks. Wu et al. [19] introduced the first continuous-time 4D LiDAR-only odometry method leveraging Doppler velocity measurements. Their method employs Gaussian process regression to estimate vehicle trajectories and correct motion distortion caused by scanning, significantly outperforming existing methods. Similarly, Hexsel et al. [7] proposed DICP, which integrates Doppler velocity measurements into the ICP framework for robust point cloud registration. By jointly optimizing Doppler and geometric objective functions, their method demonstrated improved registration accuracy and convergence speed, particularly in featureless environments such as tunnels and hallways. Yoon et al. [21] and Lisus et al. [12] developed correspondence-free methods that estimate 6-DOF velocity directly from Doppler information without requiring explicit scan matching and incorporate multiple sensors or complementary modules such as IMUs or gyroscopes for full motion estimation.
@@ II-C Summary and Limitations of 4D Radar / LiDAR Odometry
The 4D information of FMCW 4D LiDAR and 4D radar has leveraged its advantages in three main directions. First, Doppler velocity has been effectively incorporated into optimization frameworks to improve accuracy and robustness in challenging environments, such as repetitive geometric settings, foggy conditions, or a significant number of dynamic objects [3, 4, 6, 7, 11, 13, 14, 15, 19, 23]. Second, scan matching methods have been developed by refining point correspondences, using RCS information [8] or the uncertainty of point clouds [20]. Third, direct motion estimation methods [21, 12] attempt to recover full 6-DOF motion from Doppler information without relying on point correspondences.
While recent advances have shown promising results, most existing correspondence-based methods still rely solely on 3D geometric information, overlooking the rich Doppler velocity available in 4D LiDAR or radar data. Also, correspondence-free approaches estimate 6-DOF motion directly from Doppler measurements but suffer from limited rotational observability, requiring multiple LiDAR sensors or an additional gyroscope to compensate. In this work, we address these limitations by proposing a novel correspondence formulation that explicitly leverages Doppler information. Unlike prior methods, our approach associates Doppler measurements not only with translational motion but also with rotational motion from a single 4D sensor.
The remaining sections are organized as follows. In Section III, Doppler Correspondence is introduced in detail, including its geometric derivation and the rationale behind its robustness in repetitive geometric environments. Section IV presents the experimental results, demonstrating the effectiveness of the proposed method across various datasets and comparing its performance against point-to-point ICP and DICP. In Section V, the limitations of the approach are introduced, addressing scenarios where the method may face challenges. In Section VI, Doppler Correspondence is integrated with ICP to compare its performance with ICP and DICP, demonstrating its potential applicability. Finally, Section Acknowledgments provides the conclusion, summarizing the key contributions of the work and its potential for the field of 4D ranging sensor-based odometry.
@@ III-A Problem Definition
    $(\delta R,\delta t)=\underset{\delta R,\delta t}{\arg\min}\sum_{(p,q)\in\mathcal{C}}\|q-(\delta Rp\;+\;\delta t)\|_{2}.$
@@ III-D Outlier Rejection
Any pair of points that significantly deviates from these constraints is considered an outlier and is removed from the set of correspondences. By enforcing these thresholds, the impact of false correspondence estimation is mitigated.
@@
Figure 4: (a) illustrates multiple candidates due to symmetry in range and Doppler velocity, and (b) shows that the distance-based outlier rejection method filters out mismatched pairs.
@@ III-E Non-iterative Scan Matching Algorithm
The process begins by receiving two consecutive scans, $\mathcal{P}$ and $\mathcal{Q}$, and calculating $f(p_{i},v_{p,i})\text{ and }g(q_{j},v_{q,j})$. FindClosest is then performed to estimate correspondences. Unlike the conventional ICP explained in Algorithm 2, which relies on 3D-based correspondence, our method operates on 1D-based correspondence. This could reduce computational complexity, even though $f(p_{i},v_{p,i})$ and $g(q_{j},v_{q,j})$ need to be calculated. While this approach may be slightly less accurate than 3D-based methods due to less information, it offers faster computations. Furthermore, unlike the closest point correspondence methods [1, 7, 16, 17], which iteratively refine correspondences through multiple iterations, Algorithm 1 performs a direct correspondence estimation in a single step. This eliminates the iterative search process, significantly reducing the computational load, as illustrated in Fig. 1. Next section, the feasibility of Doppler Correspondence is validated by comparing the proposed method against ICP-based approaches.
@@
$\hat{R},\hat{t}\leftarrow\underset{R,t}{\arg\min}\sum_{(p,q)\in\mathcal{C}}\|q-(Rp+t)\|_{2}$ return $\hat{R},\hat{t}$
@@
      $\delta{R},\delta{t}\leftarrow\underset{R,t}{\arg\min}\sum_{(p,q)\in\mathcal{C}}\|q-(Rp+t)\|_{2}$
@@ IV-B Experiment Setup
The proposed method is evaluated by comparing its performance with established methods: point-to-point ICP and DICP [7]. DICP is a novel algorithm that integrates Doppler velocity into the Optimization in the ICP framework. It enhances scan matching by jointly optimizing both the closest point term and additional constraints derived from the Doppler velocity of each point. These methods are tested in two conditions: No Seed and With Seed. In the No Seed category, no initial pose estimate is provided, whereas the With Seed category uses the estimated pose from the previous scan pair under a constant-velocity motion model assumption [18].
@@ IV-C Evaluation Metrics
The performance is evaluated using the relative error (RE) from [24]. This metric standardizes trajectory estimation accuracy, allowing fair comparisons across different methods. The evaluation measures odometry accuracy through segment-wise error analysis. This approach provides detailed insights into the local and global performance of the trajectory estimation. By using shorter sub-trajectories, RE captures local consistency and short-term accuracy, while longer sub-trajectories emphasize global consistency and long-term accuracy. This dual capability makes RE a versatile and informative metric for trajectory analysis.
@@ IV-D Experiment Results
Figure 8: Visualization of established correspondences after a single iteration in sparse, semi-dense, and dense point clouds.
@@ IV-F Correspondence Visualization
The matching results of Doppler Correspondence and the closest point correspondence method are visualized on sparse, semi-dense, and dense point clouds. The correspondences established after the single iteration are performed, as shown in Fig. 8. With only a single iteration, Doppler Correspondence achieves more coherent and consistent matches than the closest point approach across all levels of point cloud density.
@@ V Limitations
While Doppler Correspondence, introduced in Section III, offers notable advantages in terms of computational efficiency and robustness in repetitive geometric environments, it also comes with certain limitations. Specifically, the key scenarios where the method may fail are highlighted below. We also clarify the main assumptions underlying Eqs. (6) and (7). Furthermore, potential scenarios where these assumptions may fail in practice are discussed.
@@ V-A Key Assumptions and Their Breakdown
For identification of the case where Eqs. (6) and  (7) hold, the motion of a point $p_{i}$ can be considered as a combination of translation $\Delta t\in\mathbb{R}^{3\times 1}$ and rotation $\Delta R\in SO(3)$ from the sensor between a time interval $\Delta T$. Suppose $p_{i}(T)$ is the point $p_{i}$ at time $T$. After the time interval $\Delta T$, the point moves according to:
@@
$\Delta R\;\approx\;I+\Delta T\,\omega^{\times}.$ (9)
@@
Multiplying both sides by $\Delta T$ and $\|p_{i}\|_{2}$, we obtain Eq. (6):
$(q_{j}-p_{i})\cdot p_{i}=v_{p,i}\,\Delta T\,\|p_{i}\|_{2}.$
Derivation of Eq. (7) starts from the relationship $q_{j}=\Delta Rp_{i}~{}+~{}\Delta t$, which can be expressed as:
@@
$=-v_{q,j}+(\Delta T\omega^{\times}\rho)\cdot n_{q_{j}}.$
Multiplying both sides by $\Delta T$ and $\|q_{j}\|_{2}$, we obtain:
$(q_{j}-p_{i})\cdot q_{j}=\underbrace{\Delta T~{}v_{q_{j}}\|q_{j}\|_{2}}_{\text{Doppler velocity term}}-\underbrace{\Delta T^{2}(\omega^{\times}~{}\rho)\cdot q_{j}}_{\text{rotation term}}.$
Since $\Delta T^{2}\approx 0$, the rotation term can be neglected, leading to the Eq. (7):
$(q_{j}-p_{i})\cdot q_{j}\approx v_{q_{j}}~{}\Delta T~{}\|q_{j}\|_{2}.$
Therefore, the sufficient conditions required to derive Eqs. (6) and (7) are that $\boldsymbol{\Delta T\|\omega\|_{2}\ll 1}$, as assumed in Eq. (9), and that both $\boldsymbol{\omega}$ and $\boldsymbol{\rho}$ remain constant over a short consecutive scan, as assumed in Eq. (11).
The first condition, $\Delta T\|\omega\|_{2}\ll 1$, is necessary to linearize the rotation matrix $\Delta R=\exp(\Delta T\,\omega^{\times})$. This simplification enables the relative motion between $p_{i}$ and $q_{j}$ to be approximated using linear vector operations, which is essential for establishing a direct relationship between spatial displacement and Doppler velocity.
The second condition, assuming that both $\omega$ and $\rho$ remain constant during a short interval $\Delta T$, is used in Eq. (11). This implies that the velocity of a point $p_{i}$ can be written as a time-invariant expression, ensuring that Doppler velocity measurements are consistent across a short time window. If $\omega$ or $\rho$ varies significantly, the radial velocity observed from Doppler measurements would also change, making it difficult to formulate a consistent correspondence constraint between $p_{i}$ and $q_{j}$.
Violating this condition may lead to degradation or failure of the proposed method. The following examples illustrate the cases where the assumption holds and it does not:
@@ V-A2 U-Turn (Assumption Fails)
Algorithm 3 Doppler Correspondence-Based Iterative Closest Point for 4D Ranging Sensors (Doppler 4D-ICP)
@@
      $\delta{R},\delta{t}\leftarrow\underset{R,t}{\operatorname{arg\,min}}\;(1-\alpha)\sum_{(p,q)\in\mathcal{C}_{\text{spatial}}}\rho(\|q-(Rp+t)\|_{2})+\alpha\sum_{(p,q)\in\mathcal{C}_{\text{doppler}}}\rho(\|q-(Rp+t)\|_{2})$
@@
VI Practical Use of Doppler Correspondence
Doppler Correspondence may offer versatile applications due to its simplicity. A straightforward implementation is its integration with ICP in a weighted form, referred to as Doppler 4D-ICP, as described in Algorithm 3. The process begins by computing $f(\mathcal{P})$ and $g(\mathcal{Q})$ from the point clouds and establishing Doppler Correspondence. In this setup, correspondences estimated from Doppler Correspondence remain fixed throughout the ICP iterations, while closest point correspondences are dynamically updated. This combination enables the optimization process to leverage two distinct correspondence characteristics: spatial information and Doppler information. To further enhance robustness against outliers and noise, we additionally incorporate a Huber loss function $\rho(\cdot)$ as a robust kernel during optimization.
To evaluate the effectiveness of the proposed approach, experiments are conducted using a semi-dense 4D radar dataset, Loop 1, and dense 4D LiDAR datasets, Baker-Barry Tunnel (Empty, Vehicles). The parameters are configured as follows: for the semi-dense 4D radar dataset, $\tau_{\text{spatial}}=3$, $\tau_{\text{doppler}}=10$, and $\alpha=0.7$. For the dense 4D LiDAR datasets, the parameters are set to $\tau_{\text{spatial}}=2$, $\tau_{\text{doppler}}=5$, and $\alpha=0.9$.
The experimental results are presented on the semi-dense 4D radar dataset in Table II and the dense 4D LiDAR dataset in Table III. Notably, in repetitive geometric environments, geometric ambiguity is difficult to resolve without incorporating Doppler information, leading to divergence in conventional scan-matching methods. By leveraging Doppler Correspondence, the proposed approach effectively fuses spatial and Doppler information, leading to more reliable and consistent trajectory estimation.